OpenStudy (anonymous):

simplify completely x^2+12-x/x^2-x-20 divided by 3x^2-24x+45/12x^2-48x-60

OpenStudy (earthcitizen):

\[(x^2-x+12)/((x-5) (x+4)) \div(x-3)/(4 (x+1))\]

OpenStudy (anonymous):

what does that even mean?

OpenStudy (earthcitizen):

the first equation is siplified to\[ x^2+12-x/x^2-x-20 = (x^2-x+12)/((x-5) (x+4)) \]

OpenStudy (earthcitizen):

the second equation..ii\[3x^2-24x+45/12x^2-48x-60=(x-3)/4(x+1)\]

OpenStudy (earthcitizen):

are you asked to solve completely ?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Which therefore equals\[(x^2-x+12)4(x+1)/(x-5)(x+4)(x-3)\], to intrude humbly.

OpenStudy (anonymous):

the answers are x+1/8(x-5), 8(x-5)/x+1, 4(x+1)/x-5 or x-5/4(x+1)

OpenStudy (anonymous):

PLEASE HELP :(

OpenStudy (anonymous):

i forgot the parenthese again lol

OpenStudy (anonymous):

First fraction is what you wrote correct --> x^2+12-x ???? This does not factor and I would expect it to

OpenStudy (anonymous):

no x^2+x-12 i mean that

OpenStudy (anonymous):

got it, now first fraction numerator factors to (x+4)(x-3) its denominator factors to (x-5)(x+4) and the (x+4) cancels

OpenStudy (anonymous):

for the second fraction the numerator factors to 3*(x-3)(x-5) and the denominator factors to 12*(x+1)(x-5) and the (x-5) cancels and 3/12 reduces to 1/4 So we have (x-3)/[4*(x+1)]

OpenStudy (anonymous):

Now since we are dividing, we want to flip the second and multiply across New numerator is 4*(x-3)(x+1) New denominator is (x-5)(x-3) and the (x-3) cancels You are left with [4*(x+1)]/(x-5)

OpenStudy (earthcitizen):

if the input you gave was correct then the simplified form should be\[4(x+1)/(x-5)\]

OpenStudy (anonymous):

Okay i kinda get it now thanks guys!

OpenStudy (earthcitizen):

no stress!

OpenStudy (anonymous):

do you guys get to higher levels by answering questions?